A metamer of acidic anhydride is what . Please help me

Grade:10

1 Answers

Mahima Kanawat
5 years ago
Dear student 
Let an anhydride be X-(C=O)-O-(C=O)-Y
Let X and Y be side chains here 
Metamers will form when, X and Y interchange their position  or the side chains get re arranged (same molecular formula)
WITH REGARDS 
MAHIMA 
ASKIITIANS FORUM EXPERT